# Visual Privacy Management with Generative AI for Blind and Low Vision People

### v1

@@ 1. Introduction
While these advances introduce new opportunities, they also raise concerns about visual privacy when BLV individuals leverage GenAI tools to manage, consume, and share visual content. Prior research has explored visual privacy risks in both intentionally and unintentionally sharing private content (Bennett et al., 2018; Gonzalez Penuela et al., 2022; Park, 2020; Harada et al., 2013; Adams et al., 2013; Alharbi et al., 2022; Zhang et al., 2023; Stangl et al., 2023a; Brunton and Nissenbaum, 2015; Tseng et al., 2024), human assistance versus visual assistive technologies (VATs) support (Alharbi et al., 2022; Stangl et al., 2023b), as well as in context of concern BLV users have for sharing and managing contents of their own and others with VATs (Akter et al., 2020a; Stangl et al., 2020b; Ahmed et al., 2016a). With GenAI introducing new ways of interacting visual information, it is important to understand not only in how existing visual privacy concerns may manifest differently or evolve with these emerging technologies, but also in how GenAI might actively support BLV users in safeguarding visual privacy.  [^1] Based on prior work, visual privacy refers to the safeguarding and management of sensitive visual information that could be shared or disclosed in everyday life (Stangl et al., 2023b; Ahmed et al., 2015a)
@@
Despite significant commercial and public attention towards the successes of GenAI for blind and low vision people (let, [n. d.]; see, [n. d.]; bem, [n. d.]), we know little about how blind people use GenAI tools in managing their privacy (Gurari et al., 2019; Ahmed et al., 2015b, 2016b; Voykinska et al., 2016a; Zhang et al., 2024a). Specifically, it is unclear how generative AI may be supplementing or replacing tasks traditionally supported by human assistants or VATs, and how users conceptualize and evaluate these tools in relation to their privacy concerns. Filling this gap, we explore two research questions: (RQ1:) How do blind and low vision people currently use Generative AI tools to manage visual privacy? and (RQ2:) What design opportunities do blind and low vision people envision in future Generative AI tools to support visual privacy management?
@@
To this end, we conducted semi-structured interviews with 21 individuals from the United States who are blind and have low vision to understand their current practices and design expectations for GenAI to manage visual privacy. Acknowledging the context-dependent nature of visual privacy management, through a scenario-driven inquiry, we grounded our interviews in the six different common scenarios  [^2] Six key use cases in which visual privacy can emerge: (1) Self Appearance and Impression Management, (2) Indoor Spatial Privacy, (3) Sharing Photos on Social Media, (4) Visual Privacy with Employers when Sharing Content, (5) Visual Privacy as BLV Professionals when Assessing Others’ Content, and (6) Outdoor Spatial Privacy. where privacy concerns can occur. We investigate to what extent study participants manage privacy with GenAI tools in these scenarios and then explore their expectations for how GenAI should be designed to better support visual privacy.
@@
(1) The first in-depth empirical analysis of how BLV individuals use GenAI tools to manage visual privacy. Our findings point to an emerging, yet complex, use of GenAI for visual privacy tasks, such as interpreting pregnancy tests and mammograms, despite the limitations about accuracy, to retain emotional and informational privacy in sensitive situations. Participants often emphasized values of independence and personal privacy as key factors influencing their decision to choose GenAI over VATs or human assistance.
(2) Our findings highlight the emerging use of GenAI for visual privacy management, shaped by participants’ current experiences, perceived limitations, and expectations for future design. For instance, self-appearance with GenAI is not limited to how the individual looks, but managing privacy involving bodily presentation or vulnerability to manage emotional agency. Our result also indicates privacy as institutional responsibilities where BLV professionals manage others’ content. These expectations of BLV users reflect how they negotiate trade-offs between autonomy, risk, and convenience, taking into account emotional vulnerability and privacy of others.
@@ 2.1. Use of Generative AI for Visual Accessibility
Recently, GenAI tools (e.g., ChatGPT, Google Gemini, Microsoft Copilot, and Claude) have been incorporated into a wide variety of domains, including for education (Hazzan and Erez, 2024; Kazemitabaar et al., 2023), communication (Chen et al., 2023), content creation (Inie et al., 2023), and visual interpretation (Shu et al., 2023). While much of the progress was initially for text-based applications, we are increasingly seeing a move into GenAI for visual content. [^3] Industry successes follow strong foundational research. A notable example is CLIP has shown impressive zero-shot performance (Radford et al., 2021) such as, CLIP in numerous downstream applications, ranging from object detection to 3D applications (Bangalath et al., 2022; Liang et al., 2023; Rozenberszki et al., 2022; Ni et al., 2022), and adapted for video applications (Ni et al., 2022; Wang et al., 2021; Rasheed et al., 2023). More recently, multimodal integration has advanced with models like Flamingo (Alayrac et al., 2022), BLIP-2 (Li et al., 2023b) MiniGPT-4 (Zhu et al., 2023), and LLaVA  (Liu et al., 2024) leveraging web-scale image-text data for improved multimodal chat capabilities. Some works extend LLMs for video comprehension (Maaz et al., 2023; Radford et al., 2021; Chiang et al., 2023; Li et al., 2023a; Liu et al., 2024), introducing Video-ChatGPT, a model combining a video-optimizer for enhanced understanding. Blind people are engaging with the growing ecosystem of such generative AI tools, including via accessibility technologies such as Be My Eyes and Envision AI (Vasani, [n. d.]; Alexiou, [n. d.]). In recent work on GenAI for blind users, explore accessible image generation, alt text, scene descriptions, and verification strategies. They emphasize the need for explainable, customizable, and context-aware GenAI tools to support access, creativity, and trust  (Das et al., 2024; Gonzalez et al., 2024; Chang et al., 2024; Huh et al., 2023). However, GenAI tools are rarely designed with the unique privacy needs of BLV users in mind (Adnin and Das, 2024), which could potentially result in both functional gaps and potential harms. Filling this gap, we explore how blind individuals are currently using generative AI tools for privacy management and the purposes these tools serve in their daily lives.
@@ 2.2. Visual Privacy Management with GenAI
Much of the literature on privacy management in GenAI focuses on sighted users (Sun et al., 2024; Chen et al., 2025; Wang et al., 2024; Zhou et al., 2024). Recent studies have proposed interventions to mitigate privacy risks, yet these efforts primarily address sighted users’ need, for instance, Chong et al.(Chong et al., 2024) developed a system for prompt sanitization using web-based LLMs, tackling excessive disclosure; A recurring theme across prior work is the inherent trade-off between privacy, utility, and convenience, often framed as a core challenge in GenAI. Ma et.al  (Ma et al., 2025) introduced an LLM to raise awareness of location-based privacy risks in images by detecting subtle visual cues. Similarly, CLEAR (Chen et al., 2025) provides real-time contextual risk feedback during sensitive data entry in tools like ChatGPT and Gmail, facilitating just-in-time privacy literacy. Zhou (Zhou et al., 2024) proposed Rescriber, which abstracts sensitive content before the prompt to reduce data exposure. Complementary tools such as Adanonymizer and PrivacyAsst (Zhang et al., 2024b, c) offer user-facing privacy controls for adjusting privacy-performance settings with cryptographic techniques to protect sensitive visual media. Despite these advancements, current design approaches remain limited in addressing the privacy needs of BLV users, a gap that underscores the need for inclusive GenAI privacy solutions.
@@ 2.3. BLV Individuals’ Visual Privacy Concerns and Goals
Some research revealed the visual privacy management interests of BLV individuals (Bennett et al., 2018; Gonzalez Penuela et al., 2022; Park, 2020; Harada et al., 2013; Adams et al., 2013; Ahmed et al., 2015b, 2016b; Voykinska et al., 2016a; Kaushik et al., 2023). BLV users share images and videos both to socially connect (Bennett et al., 2018; Gonzalez Penuela et al., 2022; Park, 2020) and to receive assistance in understanding their surroundings. Privacy-sensitive content includes financial and medical documents (e.g., prescriptions, pregnancy tests), identifiable personal information (e.g., addresses, faces), and images that may affect how others view them (i.e., impression management), including unflattering or awkward photographs, disorganized living spaces (Stangl et al., 2020c; Akter et al., 2020b; Sharma et al., 2023). Additional concerns arise when bystanders’ privacy is at risk or when disclosures, intentional or accidental, may impact social relationships (Stangl et al., 2020c; Akter et al., 2020b).
Pioneering visual interpretation technologies centered on human assistance, including from employees (e.g., Aira (Aira, 2020)), volunteers (e.g., Be My Eyes (BeM, 2023)), or paid crowdworkers (e.g., VizWiz (Bigham et al., 2010b)), while subsequent research showed how to reduce or eliminate human effort through automation (Wu et al., 2017, 2016; Guinness et al., 2018; Lee et al., 2022; Xie et al., 2022), eventually culminating in today’s widely used industry products, including Be My Eyes’ Be My AI app (bem, [n. d.]), Envision AI’s glasses-based technology (let, [n. d.]), and Microsoft’s Seeing AI app (see, [n. d.]). While prior research on visual privacy for blind people has primarily focused on visual interpretation services (Stangl et al., 2020c; Akter et al., 2020b), blind users’ strategies for managing privacy extend beyond this domain. Their practices intersect with broader themes in privacy research with GenAI, and contextual privacy decision-making (Nissenbaum, 2004a; Acquisti and Grossklags, 2005).In our work, we investigate how they their expectation and need can shape the design of privacy-aware GenAI.
@@ 3.1. Interview Protocol
The semi-structured interview protocol was approved by our Institutional Review Board (IRB), and can be found in Appendix A. The protocol included two sections: (a) current use of GenAI with an emphasis on visual privacy, and (b) perspectives on design improvements of GenAI in common privacy-related scenarios.
@@
In the second section, we investigated current design opportunities of GenAI by examining how it is currently used, along with participants’ perceived challenges and limitations of GenAI. To guide this investigation, we introduced a set of context-specific scenarios where privacy concerns may occur, inspired by prior work (Stangl et al., 2020a; Gurari et al., 2018; Sharma et al., 2023; Li et al., 2022; Voykinska et al., 2016b; Ahmed et al., 2015a; Feng et al., 2024; Bandukda et al., 2020). (i) self-appearance and impression management ( (Sharma et al., 2023; Brady et al., 2013; Li et al., 2022)), (ii) managing indoor spatial privacy in shared environments ( (Façanha et al., 2020; Xie et al., 2024)), (iii) sharing visual content on social media ( (Voykinska et al., 2016b; Morris et al., 2016; Ahmed et al., 2015a)), (iv) privacy in document sharing ( (Akter et al., 2020a; Feng et al., 2024; Stangl et al., 2023b)), (v) visual privacy when managing content for others (Hayes et al., 2019; Akter et al., 2022a; Wahidin et al., 2018), and (vi) outdoor spatial privacy (Bandukda et al., 2020; Jonnalagedda et al., 2014; Ahmed et al., 2015a). We designed a set of questions to (i) identify whether/ how participants currently use or would consider using GenAI for privacy management in these scenarios (see Table 1), (ii) uncover points of friction where GenAI supports or challenges established privacy expectations, and (iii) gather participants’ reasoning about potential design improvements within their privacy needs.
@@ 3.3. Data Collection & Analysis
For data collection, the first author conducted the interviews with the 21 participants remotely over Zoom between November 2024 and January 2025. The interview duration was an hour each. For data analysis, the first author led a thematic analysis (Fereday and Muir-Cochrane, 2006; Braun and Clarke, 2012) of the transcriptions alongside another with a research assistant. Two researchers independently read 20% of the interview transcripts, developed codes, and compared them until arriving at a consistent codebook. Once the codebook is ready, two researchers applied the codebook to 10% new data with inter-coder reliability calculated as 0.87 for Cohen’s Kappa (a good score as noted in (Fleiss et al., 2013)). After finalizing the codebook, the researchers divided the remaining data between them for coding and spot-checked all coded transcripts with each other to ensure ongoing consistency in coding. We followed an open coding and deductive analysis method to explore participants’ practices, challenges, and design suggestions of GenAI for visual privacy. We group 102 low-level codes into sub-themes, and further extracted main themes. We iterated this process to finally produce approximately 29 themes to interpret the results. The codebook is included in the Appendix LABEL:tab:codebook.
@@
Mamogram “I privately understand the report without asking others (P17)” Privacy Aira agents
@@
Digital & Physical Media
Email Official “I do share a lot of theoretically private information with ChatGPT, like writing letters or translating emails for official purposes.” (P12) Exploratory family, screenreader
@@ 4.2. Navigating Privacy in Digital Social Life with GenAI
Some participants shared broader concerns about the growing autonomy of GenAI tools. P15 explained how she use Seeing AI, Be my AI and chatGPT to frequently get image descriptions “I have the apps describe a photo, and I was in a dinner and I don’t share image of my child in social. I asked if there’s a child in the photo before.” P12 shared his experience of using GenAI for both image generation and privacy management-“I asked ChatGPT to turn a photo into an anime-style image of riding a rocket with sunglasses on, it was for a fun birthday post for my friend. Original photo was taken in front of my apartment, I used Be my AI to make sure my apt number was not visible.”. This highlights concerns about embedded metadata or identifiable background elements in generated images.
Similarly, P6 shared his practice for not posting certain images on social media and a similar practice in avoiding GenAI from processing images taken in certain cases. He explained “If I take a photo at a high-end restaurant on my friends’ birthday, I would not process that info in GenAI since I think I’m being surveilled, similarly I would not post it on social media. I could lose my disability benefits. I am giving AI no opportunity to take my disability benefits away from me.” This highlights a risk-aware strategy, participants navigate in response to institutional gatekeeping as well as awareness of algorithmic inference, which can invalidate one’s eligibility for public support.
Similarly, some participants mentioned using character-based GenAI tools like Kindred or Replika and charecter.AI as emotional companions. As P17 said “I use Replika with a voice I like as a private space to share personal thoughts without fear of judgment or surveillance. You know many thing I can not always do with family.” She also expressed uncertainty about potential risks, particularly whether the app retained memory of their conversations. This underscores a broader definition of privacy encompassing affective and relational dimensions where GenAI serves as a trusted confidant.

### v2

# “Before, I Asked My Mom, Now I Ask ChatGPT”: Visual Privacy Management with Generative AI for Blind and Low-Vision People

@@ 1. Introduction
While these advances introduce new opportunities, they also raise concerns about visual privacy when BLV individuals leverage GenAI tools to manage, consume, and share visual content. Prior research has explored visual privacy risks in both intentionally and unintentionally sharing private content (Bennett et al., 2018; Gonzalez Penuela et al., 2022; Park, 2020; Harada et al., 2013; Adams et al., 2013; Alharbi et al., 2022; Zhang et al., 2023; Stangl et al., 2023; Brunton and Nissenbaum, 2015; Tseng et al., 2024), human assistance versus visual assistive technologies (VATs) support (Alharbi et al., 2022; Stangl et al., 2023), as well as in context of concern BLV users have for sharing and managing contents of their own and others with VATs (Akter et al., 2020; Stangl et al., 2020b; Ahmed et al., 2016). With GenAI introducing new ways of interacting with visual information, it is important to understand not only how existing visual privacy concerns may manifest differently or evolve with these emerging technologies, but also how GenAI might actively support BLV users in safeguarding visual privacy. [^1] Based on prior work, visual privacy refers to the safeguarding and management of sensitive visual information that could be shared or disclosed in everyday life (Stangl et al., 2023; Ahmed et al., 2015)
@@
Despite significant commercial and public attention towards the successes of GenAI for blind and low vision people (let, [n. d.]; see, [n. d.]; bem, [n. d.]), we know little about how blind people use GenAI tools in managing their privacy (Gurari et al., 2019; Ahmed et al., 2015, 2016; Voykinska et al., 2016a; Zhang et al., 2024a). Specifically, it is unclear how generative AI may be supplementing or replacing tasks traditionally supported by human assistants or VATs, and how users conceptualize and evaluate these tools in relation to their privacy concerns. Filling this gap, we explore two research questions: (RQ1:) How do blind and low vision people currently use Generative AI tools to manage visual privacy? and (RQ2:) What design opportunities do blind and low vision people envision in future Generative AI tools to support visual privacy management?
@@
To this end, we conducted semi-structured interviews with 21 individuals from the United States who are blind and have low vision to understand their current practices and design expectations for GenAI to manage visual privacy. Acknowledging the context-dependent nature of visual privacy management, through a scenario-driven inquiry, we grounded our interviews in six different common scenarios [^2] Six key use cases in which visual privacy can emerge: (1) Self Appearance and Impression Management, (2) Indoor Spatial Privacy, (3) Sharing Photos on Social Media, (4) Visual Privacy with Employers when Sharing Content, (5) Visual Privacy as BLV Professionals when Assessing Others’ Content, and (6) Outdoor Spatial Privacy. where privacy concerns can occur. We investigate to what extent study participants manage privacy with GenAI tools in these scenarios and explore expectations for how GenAI should be designed to better support visual privacy.
@@
(1) The first in-depth empirical analysis of how BLV individuals use GenAI tools to manage visual privacy. Our findings point to an emerging, yet complex, use of GenAI to retain emotional and informational privacy in sensitive situations, such as interpreting pregnancy tests and mammograms, despite limitations in accuracy. Participants often emphasized values of independence and personal privacy as key factors influencing their decision to choose GenAI over VATs or human assistance.
(2) Our findings highlight the emerging use of GenAI for visual privacy management, shaped by participants’ current experiences, perceived limitations, and expectations for future design. For instance, using GenAI for self appearance is not limited to how the individual looks, but also to managing privacy involving bodily presentation or vulnerability to manage emotional agency. Our result also indicates privacy as institutional responsibilities where BLV professionals manage others’ content. These expectations of BLV users reflect how they negotiate trade-offs between autonomy, risk, and convenience, taking into account emotional vulnerability and the privacy of others.
@@ 2.1. Use of Generative AI for Visual Accessibility
Recently, GenAI tools (e.g., ChatGPT, Google Gemini, Microsoft Copilot, and Claude) have been incorporated into a wide variety of domains, including for education (Hazzan and Erez, 2024; Kazemitabaar et al., 2023), communication (Chen et al., 2023), content creation (Inie et al., 2023), and visual interpretation (Shu et al., 2023). While much of the progress was initially for text-based applications, we are increasingly seeing a move into GenAI for visual content. [^3] Industry successes follow strong foundational research. A notable example is CLIP has shown impressive zero-shot performance (Radford et al., 2021) in downstream applications, ranging from object detection to 3D applications (Bangalath et al., 2022; Liang et al., 2023; Rozenberszki et al., 2022; Ni et al., 2022), and has been adapted for video applications (Ni et al., 2022; Wang et al., 2021; Rasheed et al., 2023). More recently, multimodal integration has advanced with models like Flamingo (Alayrac et al., 2022), BLIP-2 (Li et al., 2023b) MiniGPT-4 (Zhu et al., 2023), and LLaVA  (Liu et al., 2024) leveraging web-scale image-text data for improved multimodal chat capabilities. Some works extend LLMs for video comprehension (Maaz et al., 2023; Radford et al., 2021; Chiang et al., 2023; Li et al., 2023a; Liu et al., 2024), introducing Video-ChatGPT, a model combining a video-optimizer for enhanced understanding. Blind people are engaging with the growing ecosystem of such generative AI tools, including via accessibility tools such as Be My Eyes and Envision AI (Vasani, [n. d.]; Alexiou, [n. d.]). Recent work on GenAI for blind users has explored accessible image generation, alt text, scene descriptions, and verification strategies, and has emphasized the need for explainable, customizable, and context-aware GenAI tools to support access, creativity, and trust  (Das et al., 2024; Gonzalez et al., 2024; Chang et al., 2024; Huh et al., 2023). However, GenAI tools are rarely designed with the unique privacy needs of BLV users in mind (Adnin and Das, 2024), which could potentially result in both functional gaps and potential harms. Filling this gap, we explore how blind individuals are currently using generative AI tools for privacy management and the purposes these tools serve in their daily lives.
@@ 2.2. Visual Privacy Management with GenAI
Much of the literature on privacy management in GenAI focuses on sighted users (Sun et al., 2024; Chen et al., 2025; Wang et al., 2024; Zhou et al., 2024). Recent studies have proposed interventions to mitigate privacy risks, yet these efforts primarily address sighted users’ need, for instance, Chong et al. (Chong et al., 2024) developed a system for prompt sanitization using web-based LLMs to tackle excessive disclosure. A recurring theme across prior work is the inherent trade-off between privacy, utility, and convenience, often framed as a core challenge in GenAI. Ma et al.  (Ma et al., 2025) introduced an LLM to raise awareness of location-based privacy risks in images by detecting subtle visual cues. Similarly, CLEAR (Chen et al., 2025) provides real-time contextual risk feedback during sensitive data entry in tools like ChatGPT and Gmail to facilitate just-in-time privacy literacy. Zhou (Zhou et al., 2024) proposed Rescriber, which abstracts sensitive content before the prompt to reduce data exposure. Complementary tools such as Adanonymizer and PrivacyAsst (Zhang et al., 2024b, c) offer user-facing privacy controls for adjusting privacy-performance settings with cryptographic techniques to protect sensitive visual media. Despite these advancements, current design approaches remain limited in addressing the privacy needs of BLV users, a gap that underscores the need for inclusive GenAI privacy solutions.
@@ 2.3. BLV Individuals’ Visual Privacy Concerns and Goals
Some research revealed the visual privacy management interests of BLV individuals (Bennett et al., 2018; Gonzalez Penuela et al., 2022; Park, 2020; Harada et al., 2013; Adams et al., 2013; Ahmed et al., 2015, 2016; Voykinska et al., 2016a; Kaushik et al., 2023). BLV users share images and videos both to socially connect (Bennett et al., 2018; Gonzalez Penuela et al., 2022; Park, 2020) and to receive assistance in understanding their surroundings. Privacy-sensitive content includes financial and medical documents (e.g., prescriptions, pregnancy tests), identifiable personal information (e.g., addresses, faces), and images that may affect how others view them (i.e., impression management), including unflattering or awkward photographs, disorganized living spaces (Stangl et al., 2020b; Akter et al., 2020; Sharma et al., 2023). Additional concerns arise when bystanders’ privacy is at risk or when disclosures, intentional or accidental, may impact social relationships (Stangl et al., 2020b; Akter et al., 2020).
Pioneering visual interpretation technologies centered on human assistance, including from employees (e.g., Aira (Aira, 2020)), volunteers (e.g., Be My Eyes (BeM, 2023)), or paid crowdworkers (e.g., VizWiz (Bigham et al., 2010b)), while subsequent research showed how to reduce or eliminate human effort through automation (Wu et al., 2017, 2016; Guinness et al., 2018; Lee et al., 2022; Xie et al., 2022), eventually culminating in today’s widely used industry products, including Be My Eyes’ Be My AI app (bem, [n. d.]), Envision AI’s glasses-based technology (let, [n. d.]), and Microsoft’s Seeing AI app (see, [n. d.]). While prior research on visual privacy for blind people has primarily focused on visual interpretation services (Stangl et al., 2020b; Akter et al., 2020), blind users’ strategies for managing privacy extend beyond this domain. Their practices intersect with broader themes in privacy research with GenAI, and contextual privacy decision-making (Nissenbaum, 2004; Acquisti and Grossklags, 2005). In our work, we investigate how they their expectation and need can shape the design of privacy-aware GenAI.
@@ 3.1. Interview Protocol
The semi-structured interview protocol was approved by the Institutional Review Board (IRB) at Penn State, and can be found in Appendix A. The protocol included two sections: (a) current use of GenAI with an emphasis on visual privacy, and (b) perspectives on design improvements of GenAI in common privacy-related scenarios.
@@
In the second section, we investigated current design opportunities of GenAI by examining how it is currently used, along with participants’ perceived challenges and limitations of GenAI. To guide this investigation, we introduced a set of context-specific scenarios where privacy concerns may occur, inspired by prior work (Stangl et al., 2020a; Gurari et al., 2018; Sharma et al., 2023; Li et al., 2022; Voykinska et al., 2016b; Ahmed et al., 2015; Feng et al., 2024; Bandukda et al., 2020). (i) self-appearance and impression management ( (Sharma et al., 2023; Brady et al., 2013; Li et al., 2022)), (ii) managing indoor spatial privacy in shared environments ( (Façanha et al., 2020; Xie et al., 2024)), (iii) sharing visual content on social media ( (Voykinska et al., 2016b; Morris et al., 2016; Ahmed et al., 2015)), (iv) privacy in document sharing ( (Akter et al., 2020; Feng et al., 2024; Stangl et al., 2023)), (v) visual privacy when managing content for others (Hayes et al., 2019; Akter et al., 2022; Wahidin et al., 2018), and (vi) outdoor spatial privacy (Bandukda et al., 2020; Jonnalagedda et al., 2014; Ahmed et al., 2015). We chose these scenarios based on (a) relevancy of the scenario to visual privacy with GenAI ; (b) gap identified in prior work in visual privacy; and (c) diversity of context representing BLV users. Our goal is to fill the gap in exploring the limitations and design opportunities for these use-cases for GenAI. We designed a set of questions to (i) identify whether/ how participants currently use or would consider using GenAI for privacy management in these scenarios (see Table 1), (ii) uncover points of friction where GenAI supports or challenges established privacy expectations, and (iii) gather participants’ reasoning about potential design improvements within their privacy needs.
@@ 3.3. Data Collection & Analysis
For data collection, the first author conducted the interviews with the 21 participants remotely over Zoom between November 2024 and January 2025. The interview duration was an hour each. For data analysis, the first author led a thematic analysis (Fereday and Muir-Cochrane, 2006; Braun and Clarke, 2012) of the transcriptions alongside another with a research assistant. Two researchers independently read 20% of the interview transcripts, developed codes, and compared them until arriving at a consistent codebook. Once the codebook is ready, two researchers applied the codebook to 10% new data with inter-coder reliability calculated as 0.87 for Cohen’s Kappa (a good score as noted in (Fleiss et al., 2013)). After finalizing the codebook, the researchers divided the remaining data between them for coding and spot-checked all coded transcripts with each other to ensure ongoing consistency in coding. We followed an open coding and deductive analysis method to explore participants’ practices, challenges, and design suggestions of GenAI for visual privacy. We group 102 low-level codes into sub-themes, and further extracted main themes. We iterated this process to finally produce approximately 29 themes to interpret the results.
@@
Mammogram “I privately understand the report without asking others (P17)” Privacy Aira agents
@@
Digitized Physical Media
Official email “I do share a lot of theoretically private information with ChatGPT, like writing letters or translating emails for official purposes.” (P12) Exploratory family, screenreader
@@ 4.2. Navigating Privacy in Digital Social Life with GenAI
Some participants shared broader concerns about the growing autonomy of GenAI tools. P15 explained how she use Seeing AI, Be my AI, and chatGPT to frequently get image descriptions “I have the apps describe a photo, and I was in a dinner and I don’t share images of my child in social. I asked if there’s a child in the photo before.” P12 shared his experience of using GenAI for both image generation and privacy management-“I asked ChatGPT to turn a photo into an anime-style image of riding a rocket with sunglasses on, it was for a fun birthday post for my friend. Original photo was taken in front of my apartment. I used Be my AI to make sure my apt number was not visible.”. This highlights concerns about embedded metadata or identifiable background elements in generated images.
Similarly, P6 shared his practice for not posting certain images on social media and a similar practice in avoiding GenAI from processing images taken in certain cases. He explained “If I take a photo at a high-end restaurant on my friends’ birthday, I would not process that info in GenAI since I think I am being surveilled, similarly, I would not post it on social media. I could lose my disability benefits. I am giving AI no opportunity to take my disability benefits away from me.” This highlights a risk-aware strategy, participants navigate in response to institutional gate-keeping as well as awareness of algorithmic inference, which can invalidate one’s eligibility for public support.
Similarly, some participants mentioned using character-based GenAI tools like Kindred or Replika and character.AI as emotional companions. As P17 said “I use Replika with a voice I like as a private space to share personal thoughts without fear of judgment or surveillance. You know many thing I can not always do with family.” She also expressed uncertainty about potential risks, particularly whether the app retained memory of their conversations. This underscores a broader definition of privacy encompassing affective and relational dimensions where GenAI serves as a trusted confidant.